OpenStudy (anonymous):

The driver of a car traveling at 10.0 m/s along a straight road depresses the accelerator and uniformly increases her speed at a rate of 2.20 m/s2. How fast will the car be moving as it passes a parked police cruiser 90 m away?

OpenStudy (stormfire1):

(Assuming you are talking 90 meters and not miles) Start with the known values: V0 = 10m/s X0 = 0m a = 2.2m/s^2 Xf = 90m Now just use the kinematic formula which allows you to solve for Vf: \[Vf ^{2}=Vi ^{2}+2a(Xf-Xi)\] So it's \[Vf^2 = (10m/s^2)^2 + 2 (2.2m/s^2)(90m) = 496 m/s\]\[V = \sqrt{496}=22.27 m/s\]

OpenStudy (anonymous):

Thanks a lot! that helped me :)